A business uses linear depreciation for its machines. The depreciation time is set to five years. The machines were purchased at the entrance to Year 1 for NOK 3.000.000 including VAT ( Norwegian mva). Residual Value at the expiration of the service life was estimated at NOK 400.000 without VAT(Norwegian mva). After four years the machine was sold. The loss of sales was NOK 30.000.
What was the sales total including VAT ( Norwegian mva)?
1
Expert's answer
2019-10-31T10:34:50-0400

890.000

(3.000.000-400.000)/5=520.000 - depreciation per year

3.000.000-520.000*4=920.000 - price at the end of the forth year

920.000-30.000=890.000